OpenStudy (anonymous):

write this fraction as a mixed number or whole number in lowest terms 35/11= 3 2/11 42/18= 2 1/6 100/12 = 8 1/3 RIGHT? write this fraction as a mixed number or whole number in lowest terms 35/11= 3 2/11 42/18= 2 1/6 100/12 = 8 1/3 RIGHT? @Mathematics

OpenStudy (turingtest):

42/18=14/6=7/3=2 1/3 the others are correct